I have a question that is driving my wife and I crazy. Can you help us figure it out. It goes something like this:
Leah is 5 yrs older than Sue. John is 6 yrs older than Leah. Their ages total equals 41. How old is Sue? What is the equation for this problem? Thank you.

Answer by longjonsilver(2297)

this has 2 unknowns, so you will need 2 equations.

First define a couple of things:
Let x = Leah's age now
Let y = Sue's age now
eqn1: x = y+5 --> x-y = 5
eqn2: x+y = 41

Adding these 2 equations together gives 2x = 46, so x (Leah's age) = 23. Hence, Sue must be 18.

jon.